OpenStudy (anonymous):

What is LCM of p^2 and p^2+7p+10

OpenStudy (anonymous):

p^4+7p^3+10p^2

OpenStudy (anonymous):

that is not correct

OpenStudy (anonymous):

Break down both into prime factors so to speak which is p^2 and (p+2)(p+5) There is no highest common factor other than 1 Therefore they have to be multiplied to find LCM Which is p^2 ( p^2+7p+10) giving p^4+7p^3+10p^2 Think about finding the LCM of 12 and 35 = (3.2^2)and (3.7) the HCF is 1 so we multiply to find LCM which is 35 x 12 = 420 The same applies to algebra take a different problem LCM of x^2 +3x +2 and x^2 +6x +5 Factored we get x^2 +3x +2 = (x+1)(x+2) x^2 +6x +5 = (x+1)(x+5) (x+1) is the HCF of both expressions so that is only needed once. Therefore the LCM is (x+1)(x+2)(x+5)